OpenStudy (anonymous):

Functions and inverses If ƒ(x ) = 3x + 1 and ƒ ^-1 = x-1/3 then ƒ(2) = 1/3 1 7

OpenStudy (anonymous):

Any help would be great.

OpenStudy (freckles):

f(2) means you just replace the x's in the function named f above

OpenStudy (freckles):

you have f(x)=3x+1 you want to find f(2) replace the x's with 2

OpenStudy (anonymous):

7 or 1/3

OpenStudy (anonymous):

?

OpenStudy (freckles):

well you are to find f(2) not f^(-1)(2)

OpenStudy (freckles):

so what do you think?

OpenStudy (anonymous):

So 7

OpenStudy (freckles):

yes

OpenStudy (anonymous):

3*2+1= 7

OpenStudy (anonymous):

Thanks.

OpenStudy (freckles):

np

OpenStudy (anonymous):

Can I check 1 more? If ƒ(x ) = 3x + 1 and ƒ ^-1 = x-1/3 then ƒ ^-1(7) = 22 2 -2

OpenStudy (anonymous):

7-1=6/3=2

OpenStudy (freckles):

is that really (x-1)/3 or is it x-1/3 judging by what f is I believe f^-1(x)=(x-1)/3 and if so f^-1(7)=(7-1)/3=6/3=2 is right

OpenStudy (anonymous):

x-1 is over 3

OpenStudy (anonymous):

or divided by 3

OpenStudy (anonymous):

which ever terminology you prefer